# Enhancing Event Extraction from Short Stories through Contextualized Prompts

### v1

@@ 5.2 Data Augmentation
The gold standard data is necessary to create any Natural Language Processing (NLP) application. The annotation process has become more difficult as NLP has attempted to handle increasingly complex linguistic phenomena or a variety of labelling and classification tasks, yet the requirement for and usefulness of large labelled datasets persist [sun2017revisiting, human-annotation]. Manual annotating linguistic resources for training and testing is expensive, especially for specialized text fields. For the same reason, we tried to generate the labels for event extraction and classification automatically from EVENT PROMPT. The approach incorporates the following steps:
@@
    Train some models on the manually annotated data to benchmark it.
@@
    Utilize the best-performing trained models to synthetically generate labels for the 800 unlabelled short stories, which can be further enhanced by utilizing human expertise.
@@ 5.3 Ablation Study
    Effect of Learning Rate: We performed an ablation study to see the effect of different learning rates on the EVENT PROMPT model represented in Table 10. The model achieved its best performance with a learning rate of 4e-5, achieving a macro F1 score of 74.74%.
@@
    Training with 120 synthetic short stories: just discussed in earlier subsection
@@
    Training with 500 synthetic short stories: just discussed in earlier subsection
  4. 4.
    What to include????
@@ 6 Error Analysis
This section describes the errors while predicting realis events using the best-performing EVENT PROMPT model. We observed that most errors stemmed from incorrectly classifying unreal events as real ones. In the sentence, “He dreamt that he was flying in the air.” the word “flying” was predicted as a realis event by the BiLSTM+Document context model. However, “flying” is happening in someones dream, which is an unreal event. Similarly, in the sentence, “While passing by, the two bulls Cleo and Gabby would stop seeing the garbage.” the word “seeing” was classified as realis event. However, it is clear that the sentence is a generic statement and cannot be treated as a realis event according to the aspects we mentioned in LABEL:fig:realis. Furthermore, in the zero-shot setting of the prompt-based learning method, the GPT-4 was even predicting proper nouns, such as the names of persons or places as events. However, after providing 30 examples of realis events as feedback (30 shots), most of the error was found to be coming from predicting the events that might happen in the future. For instance, in the sentence, “You will be surprised to know that Hindi is now being taught in schools.” the word “surprised” was wrongly predicted as a realis event. Apart from the error, it is noteworthy that the prompt-based approach has demonstrated its ability to detect events even within complex sentences successfully. For instance, in the sentence, “Sunny was surprised to see the police at his farm.”. The GPT-4 has predicted only the word “surprised” as the event while neglecting the word “see”. This was in accordance with the annotation guideline.
6.1 Limitation and Challenges
@@ 7 Conclusions
In this work, we proposed a prompt-based approach for event trigger detection and classification in short stories using contextualized prompts. We first compared traditional feature extraction models with the prompt-based approaches. We tried to address problems faced when working with event extraction, like the unavailability of good-quality annotated datasets and data imbalances. We have used a pre-trained language model built on the standard transformer, BART. The results show that the proposed method outperforms the baseline models. Since the target audience for short stories is children, we observed an imbalance in the number of events in the dataset (e.g. less number of events for CONFLICT class and more for COMMUNICATION class). It was also observed that the heavily imbalanced classes improved through contextualized prompts. We got a performance improvement from 65.2 to 69.4% in F1-score for CONFLICT category, which had a significantly less number of instances throughout the dataset. We have also shown how we could integrate the Event Extraction framework by first detecting the events followed by Event Classification. The best-performing model (EVENT PROMPT) was used to expand the dataset and annotate 800 stories synthetically, making a complete dataset of 1000 short stories. This synthetically generated dataset was validated with extensive experiments. Prompt-based methods are, therefore, found to be efficient for small and imbalanced datasets. Overall, contextualized prompts are a powerful tool for enhancing the performance of event extraction models by providing them with the necessary context and guidance to extract events accurately from text and classify them. The domain of Prompt-Based learning is growing and is flexible enough to be used in most information retrieval tasks. There are a lot of scopes available in argument mining and other subtopics.

### v2

@@ 6.1 Data Augmentation
The gold standard data is necessary to create any Natural Language Processing (NLP) application. The annotation process has become more difficult as NLP has attempted to handle increasingly complex linguistic phenomena or a variety of labelling and classification tasks, yet the requirement for and usefulness of large labelled datasets persist [Sun et al., 2017, Liu, 2023]. Manual annotating linguistic resources for training and testing is expensive, especially for specialized text fields. For the same reason, we tried to generate the labels for event extraction and classification automatically from EVENT PROMPT. The approach incorporates the following steps:
@@
    Train EVENT PROMPT models on the manually annotated data.
@@
    Utilize EVENT PROMPT model to synthetically generate labels for the 800 unlabelled short stories, which can be further enhanced by utilizing human expertise.
@@ 6.2 Ablation Study
    Effect of Learning Rate: We conducted an ablation study to evaluate the impact of various learning rates on the performance of the EVENT PROMPT model. The results, presented in Table 10, indicate that the model achieved its highest performance with a learning rate of 4e-5. At this optimal learning rate, the model attained a macro F1 score of 74.74%, highlighting the sensitivity of the model’s performance to changes in the learning rate parameter. This finding underscores the importance of carefully tuning hyperparameters to enhance model effectiveness.
@@
    Training with 120 synthetic short stories: Training the EVENT PROMPT model with 120 synthetic short stories resulted in degraded performance. The impact of using Silver standard data is clearly reflected in the model’s performance. This degradation suggests that the quality of the training data plays a crucial role in the effectiveness of the model. High-quality, authentic data may be necessary to achieve optimal results, as synthetic or lower-quality data can negatively affect the model’s ability to generalize and perform well on real-world tasks.
@@
    Training with 500 synthetic short stories: Furthermore, the EVENT PROMPT model was trained with 500 synthetically labelled short stories. The performance results, shown in Table 8 and Table 9, indicate that the degradation in performance caused by the Silver standard data can be mitigated by increasing the volume of such data. This suggests that while Silver standard data may initially lead to a decrease in model performance, a larger quantity of this data can help compensate for its lower quality, ultimately enhancing the model’s overall effectiveness. This finding underscores the potential of using extensive synthetically labeled datasets to improve model performance when high-quality labeled data is limited.
@@ 7 Error Analysis
This section describes the errors encountered while predicting realis events using the best-performing EVENT PROMPT model. The errors can be categorized into two main types based on the subtasks: event detection and event classification.
For the task of event detection, it was observed that most errors stemmed from incorrectly classifying unreal events as real ones. This might be due to subtle contextual cues or ambiguities in the training data. Although the performance was fairly good, the thin line between realis and non-realis events sometimes creates confusion. Improving the model’s ability to detect these nuances may require more sophisticated feature extraction or additional training on more diverse and representative datasets.
In the context of event classification, the class with the least performance is OTHERS. This is due to the presence of an almost negligible amount of this type of event in the dataset. The model didn’t have enough samples to train on. Additionally, the misclassification between the GENERAL-ACTIVITY and OTHERS categories was higher because the GENERAL-ACTIVITY class has the widest scope of events, while the OTHERS category acts as an outlier class.
@@ 8 Conclusions
In this work, we proposed a prompt-based approach for event trigger detection and classification in short stories using contextualized prompts. We first compared traditional feature extraction models with the prompt-based approaches. We tried to address problems faced when working with event extraction, like the unavailability of good-quality annotated datasets and data imbalances. We have used a pre-trained language model built on the standard transformer, BART. The results show that the proposed method outperforms the baseline models. Since the target audience for short stories is children, we observed an imbalance in the number of events in the dataset (e.g. less number of events for CONFLICT class and more for COMMUNICATION class). It was also observed that the heavily imbalanced classes improved through contextualized prompts. We got a performance improvement from 65.2 to 69.4% in F1-score for CONFLICT category, which had a significantly less number of instances throughout the dataset. The best-performing model EVENT PROMPT was used to expand the dataset and annotate 800 stories synthetically, making a complete dataset of 1000 short stories. This synthetically generated dataset was validated with extensive experiments. Prompt-based methods are, therefore, found to be efficient for small and imbalanced datasets. Overall, contextualized prompts are a powerful tool for enhancing the performance of event extraction models by providing them with the necessary context and guidance to extract events accurately from text and classify them. The domain of Prompt-Based learning is growing and is flexible enough to be used in most information retrieval tasks. There are a lot of scopes available in argument mining and other subtopics.